Calculate the mass of manganese (IV) oxide that can be synthesized from 15.00 grams of potassium iodide.


1
Expert's answer
2021-10-15T12:31:54-0400

6KI(aq) +4H2O(l)+2KMnO4(aq)"\to" 3l2(s) +2 MnO2(s)+8KOH(aq)


REM of KI=166.00 g/mol

Moles of KI=15/ 166

= 0.09036moles

Mole ratio= KI:MnO2 = 6:2 = 3:1

Moles of MnO2= 0.09036/3

=0.03012moles

REM of MnO2= 86.99 g/mol

Mass of MnO2= 0.03012× 86.97

=2.6195364g